### v1

@@ Abstract
Human pose estimation (HPE) from Radio Frequency vision (RF-vision) performs human sensing using RF signals that penetrate obstacles without revealing privacy (e.g., facial information). Recently, mmWave radar has emerged as a promising RF-vision sensor, providing radar point clouds by processing RF signals. However, the mmWave radar has a limited resolution with severe noise, leading to inaccurate and inconsistent human pose estimation. This work proposes mmDiff, a novel diffusion-based pose estimator tailored for noisy radar data. Our approach aims to provide reliable guidance as conditions to diffusion models. Two key challenges are addressed by mmDiff: (1) miss-detection of parts of human bodies, which is addressed by a module that isolates feature extraction from different body parts, and (2) signal inconsistency due to environmental interference, which is tackled by incorporating prior knowledge of body structure and motion. Several modules are designed to achieve these goals, whose features work as the conditions for the subsequent diffusion model, eliminating the miss-detection and instability of HPE based on RF-vision. Extensive experiments demonstrate that mmDiff outperforms existing methods significantly, achieving state-of-the-art performances on public datasets.
@@ 1 Introduction
Overcoming limitations of camera-based HPE, Radio-Frequency vision (RF-vision) has attracted surging attention for human sensing. The emerging mmWave radar technology presents a promising and feasible solution for HPE due to its privacy preservation, portability, and energy efficiency [43]. Operated at the frequencies of 30-300 GHz [49], commercial mmWave radars transmit and receive RF signals that penetrate human targets and occlusions. Radar point clouds (PCs) are further extracted as the salient target detection via monitoring signals’ characteristic changes [33, 46]. Therefore, the extracted PCs are more robust to adverse environments [51] and reveal little privacy [47], which inspires accurate and privacy-preserving HPE [23, 47, 2, 46]. However, due to the bandwidth and hardware limitations [29], radar PCs are sparse with limited geometric information obtained [2], leading to huge difficulties in achieving HPE. The sparse PCs are noisy in two aspects: (1) mmWave radar has a lower spatial resolution, leading to PC dispersion throughout the target area accompanied by ghost points caused by multi-path effect [40, 29]; (2) signal’s specular reflection and interference [6, 5] further cause inconsistent sensing data, leading to occasional miss-detection of human parts.
Figure 1: Left: challenges of mmWave PCs. Right: the performance of existing SOTA (P4Transformer [13]) compared to ours. PC’s sparsity and dispersion cause inaccurate spline and shoulder. Inconsistent PCs with occasional miss-detection further cause size variance and pose vibration. mmDiff proposes diffusion-based pose estimation with enhanced accuracy and stability.
@@ 3 Methodology
Figure 2: mmDiff proposes diffusion-based pose estimation with a conditional diffusion model, using mmWave radar information as conditions. $k\in[0..K]$ denotes the diffusion step. Four modules are proposed to extract more reliable guidance, addressing PCs’ noise and inconsistency respectively: GRC and LRC first extract more robust global-local radar features, $C^{glo}$ and $C^{loc}$; SLC and TMC then extract consistent human structure and motion patterns, $C^{tem}$ and $C^{lim}$.
We address the task of 3D human pose estimation (HPE) with noisy and sparse point clouds (PCs) from mmWave radar. At each time frame $t\geq 0$, given the input 6-dimensional radar point cloud $R_{t}\in\mathbb{R}^{N\times 6}$, where $N$ denotes the number of detected points, our task is to estimate the ground-truth 17-joint human pose $H_{t}=\{h_{t}^{1},h_{t}^{2},...,h_{t}^{17}\}\in\mathbb{R}^{17\times 3}$. Apart from the Cartesian coordinate $\{x,y,z\}$, each detected radar point also contains three attributes $\{v,E,A\}$, where $v$ denotes velocity, $E$ denotes energy, and $A$ denotes amplitude.
@@ 3.1 Diffusion-based Human Pose Estimation
Diffusion models [16, 38] are a category of probabilistic generative models popular in various generative tasks, e.g., image generation. Given a noisy image from a noisy distribution, diffusion models can generate realistic image samples that match the natural image distribution, through iterative noise removal [25, 19]. Extended to HPE, diffusion models can estimate the distribution of reasonable human poses for realistic pose generation. Particularly with noisy radar modality, miss-detected body joints can be accurately estimated by inferring from the detected ones, and inaccurate joints causing twisted human structures are potentially refined.
@@ 3.2 Global-local Radar Context
Existing mmWave HPE paradigm applies encoder-decoder architectures to encode the radar point clouds (PCs), $R\in\mathbb{R}^{N\times 6}$, into latent feature representations and decode them into human poses. Anchor-based methods [13, 50] are common options for PC encoders, where point anchors are first sampled from the Farthest Point Sampling algorithm and nearby PCs are extracted as the anchor features. As a result, the holistic radar PCs are encoded into PC features $F^{r}\in\mathbb{R}^{P\times 1024}$, where $P$ indicates the anchor number. To decode the PC features, a Multi-Layer Perceptron (MLP) is commonly applied as a dimension-reducing projection to generate the joint feature $F^{j}\in\mathbb{R}^{17\times 1024}$, which is further decoded into human poses $\tilde{H}\in\mathbb{R}^{17\times 3}$ by another MLP. However, occasional radar miss-detection causes uncertainty within the PC features $F^{r}$, while MLP-based projection can hardly identify miss-detected joints. Additionally, radar low-resolution PCs further impose noises. As a result, the estimated human joints are generally coarse.
@@
LRC further performs local point-to-point self-attention to extract higher-resolution features from local PCs near body joints. To select local PCs, existing methods [46] utilize static joint anchors from coarse-estimated human poses $\tilde{H}$. However, the local joint features should dynamically reflect the joints’ errors at different diffusion steps. Therefore, we propose dynamic joint anchors from intermediate diffusion poses $\hat{H}_{k}$ for PC selection. With $i\in[1,..,17]$ and each joint $\hat{h}_{k}^{i}$ from the $\hat{H}_{k}$ as an anchor, the $\bar{K}$-nearest-neighbors ($\bar{K}$) algorithm is applied to select $\bar{K}$ nearest points as local PCs, $\bar{R}^{i}_{k}\in\mathbb{R}^{\bar{K}\times 6}$. Each $\bar{R}^{i}_{k}$ is first encoded by a shared MLP $g^{l}$ into a $\mathbb{R}^{\bar{K}\times 64}$ embedding, and then fed into a shared small-scale local transformer $\Phi^{l}$ for point-to-point self-attention. Finally, average pooling is performed to generate $\mathbb{R}^{64}$ embeddings, which are further aggregated (concatenated) for every joint anchor $\tilde{h}^{i}$ as the conditional embedding:
@@ 3.4 Overall Learning Objective
Then, in phase two, the remaining three conditional modules and the diffusion model are trained together. The extracted $F^{j}$ and $\tilde{H}$ serve as the input of structural-motion consistency modules, and $\tilde{H}$ initialize $\hat{H}_{K}$ for the reverse diffusion process. The learning objective is minimizing $\mathcal{L}_{\text{diff}}$, which is the diffusion learning objective following DDPM [16]. To ensure accurate limb-length estimation, a $\mathcal{L}_{1}$ limb regression loss is further designed and integrated:
@@ 4 Experiments
Section 4.1 presents our experiment setting. Section 4.2 demonstrates the overall performance of mmDiff on two public datasets: mmBody [6] and mmFi [47]. Section 4.3 further performs ablation studies to validate the designed modules, followed by analytics of pose stability and module efficiency.
@@ 4.1 Experiment Setup
mmBody [47] studies the robustness of human sensing with various sensors: RGB, Depth, and mmWave radar. Human skeletons are annotated by the MoCap system. Models are set to train on data collected from 2 standard scenes (Lab1 and Lab2), and tested on 3 basic scenes and 4 adverse scenes (including unseen subjects). Meanwhile, mm-Fi [47] is a larger-scale dataset consisting of 25 activities and 40 subjects using a low-cost mmWave radar (sparser PCs). The skeleton annotations are rather unstable compared to the MoCap annotations, as obtained by RGB using the pretrained HRNet-w48 [39]. We test our methods based on Protocol 1, with all daily activities. Three data-splitting methods are used with a train-test split ratio of 4:1: random split, cross-subject split, and cross-environment split.
@@
The GCN encoders, GCN blocks, and GCN decoders are designed following [52] and [15]. All MLPs are implemented as (LayerNorm, Linear, Dropout, ReLU, and Linear), and the temporal 1D-convolution extractor $g_{2}^{tem}$ is implemented as (Conv1D(k=3), Dropout, ReLU, Conv1D(k=3), MaxPool).
Our methods are trained for 100 epochs with a fixed batch size of 1024. The Adam algorithm [20] algorithm is used for optimization, with the learning rate set as $52e-$ 5, the gradient clip set to 1.0, and the momentum set to 0.9. The length of the forward/reverse diffusion process is set as $K=25$ with the constant $\beta$ sampling of 0.001. An average of 5 hypotheses is recorded for fair comparison with non-diffusion methods. We choose Point4D [13] as the GRC’s PC encoder for mmBody and PointTransformer [50] for mm-Fi. We set $\bar{K}=50$ for $\bar{K}$ algorithm for LRC, but due to insufficient radar points ($N<100$), the LRC module is neglected for mm-Fi. We further set $\Delta t=8$ for TMC and $\lambda=5$ for SLC. The hyper-parameters are obtained empirically, more precise hyper-parameter tuning tricks such as the Bayesian optimization could lead to better results.
4.1.3 Compared Methods.
(1) RGB, Depth, and RGB(D) [6] are benchmarks using different modalities. (2) mmWave methods: P4Transformer [13] for mmBody contains Point4D Convolution as the PC encoder and a transformer for self-attention; PointTransformer [50] is designed to handle the sparse PCs in mm-Fi. mmMesh [46] is implemented as an extra mmWave HPE baseline on mmBody. (3) SOTA camera-based HPE methods: camera-based 3D HPE methods generally perform pose lifting from 2D poses to 3D poses. To modify them to perform radar-based HPE, we train the models to perform pose refinements, from coarse poses $\tilde{H}$ to clean 3D poses. PoseFormer [53] is the transformer-based method with temporal-spatial attention. DiffPose [15] is the diffusion-based method using SOTA graph-based GraFormer [52] as backbones.
4.1.4 Evaluation Metric.
Two evaluation metrics are adopted following [18]: (1) Mean Per Joint Position Error (MPJPE (mm)): the average joint error between ground truth and prediction (after pelvis alignment), indicating keypoint positional correctness; and (2) Procrustes Analysis MPJPE (PA-MPJPE (mm)): procrustes methods (translation, rotation, and scaling) are performed before error calculation, indicating the overall human pose quality.
@@
PointTransformer [50]
$73.09_{\pm 2.70}$ $55.60_{\pm 1.40}$ $75.96_{\pm 6.90}$ $58.70_{\pm 4.30}$ $88.28_{\pm 4.50}$ $68.79_{\pm 2.79}$
Diffpose* [15]
$73.44_{\pm 0.29}$ $56.83_{\pm 0.25}$ $70.31_{\pm 0.27}$ $54.12_{\pm 0.31}$ $86.35_{\pm 0.06}$ $66.87_{\pm 0.17}$
mmDiff(G)*
$68.62_{\pm 0.06}$ $53.11_{\pm 0.05}$ $68.46_{\pm 0.06}$ $52.55_{\pm 0.05}$ $85.63_{\pm 0.53}$ $66.43_{\pm 0.28}$
mmDiff(G,S)*
$65.72_{\pm 0.08}$ $50.72_{\pm 0.01}$ $67.18_{\pm 0.18}$ $51.85_{\pm 0.05}$ $83.39_{\pm 0.17}$ $64.61_{\pm 0.40}$
mmDiff(G,S,T)*
$\textbf{65.26}_{\pm\textbf{0.11}}$ $\textbf{50.35}_{\pm\textbf{0.09}}$ $\textbf{65.62}_{\pm\textbf{0.24}}$ $\textbf{50.23}_{\pm\textbf{0.24}}$ $\textbf{82.73}_{\pm\textbf{0.62}}$ $\textbf{63.87}_{\pm\textbf{0.26}}$
Figure 3: Qualitative visualization of the estimated poses on mmBody dataset. mmDiff demonstrates higher keypoint accuracy.
@@ 4.2 Overall Result
Figure 4: (a) qualitatively shows the pose motion stability on mm-Fi, by plotting 5 consecutive frames of estimated poses. mmDiff shows more consistent motion patterns (zoom in for details). (b) shows the motion energy levels on mmBody, calculated by the Average Keypoint Velocity (AKV). Lower AKV indicates better stability.
Table 3: Ablation studies of proposed modules on mmBody.
@@ 4.3 Analytics
In Figure 4 (a), we observe a mixing of incorrect skeletons within the correct skeleton timeline with PointTransformer [50], such as chest-expanding and hand-waving. With mmDiff, smooth motion patterns are ensured based on human behavioral prior knowledge. Additionally, our proposed mmDiff can mitigate the uncertain locations of legs and arms (caused by low resolution), demonstrating enhanced pose accuracy. Furthermore, for throwing and kicking actions, our method is more stable compared to RGB-extracted ground truth, as camera-based HPE suffers from self-occlusion. In Figure 4 (b), we further apply the Average Keypoint Velocity (AKV) to quantitatively measure the pose stability. AKV is defined as $E_{i=[0..17]}(||J_{t}^{i}-J^{i}_{t-1}||^{2})$, which measures the average inter-frame joint moving distance, indicating the motion energy level and pose stability. The proposed mmDiff demonstrates enhanced pose stability by minimizing joint vibration and avoiding abrupt pose changes.
@@
To validate the effectiveness of the limb-length consistency, we compare mmDiff w/ and w/o the module ($C^{lim}$) and plot the histograms that indicate the limb-length distribution in Figure 5. The ground truth limb length remains constant for all limbs. We observe both reduced limb-length error (in the estimated arms, legs, and spline) and variance (in forearms and lower legs), indicating enhanced pose accuracy and structural stability. We argue that more accurate limb-length can lead to more accurate pose estimation. Qualitatively in figure 4, the variant height and arm’s length for chest expanding and hand raising are mitigated with mmDiff.
@@
We provide the efficiency analysis in Table 5, where our proposed modules require substantially little computational resources during phase-two diffusion training. We observe that all designed modules have a latency of less than 2ms, demonstrating outstanding efficiency in supporting the iterative diffusion process. Meanwhile, as reflected by the module’s parameters, the model size of our designed modules is relatively small compared to the P4Transformer [13], leading to more stable training convergence. Moreover, the efficient module design requires minimum computation complexity, as reflected by GFLOPs, demonstrating the potential for applications like robotics, the Internet of Things, and edge computing.

### v2

@@ Abstract
Human pose estimation (HPE) from Radio Frequency vision (RF-vision) performs human sensing using RF signals that penetrate obstacles without revealing privacy (e.g., facial information). Recently, mmWave radar has emerged as a promising RF-vision sensor, providing radar point clouds by processing RF signals. However, the mmWave radar has a limited resolution with severe noise, leading to inaccurate and inconsistent human pose estimation. This work proposes mmDiff, a novel diffusion-based pose estimator tailored for noisy radar data. Our approach aims to provide reliable guidance as conditions to diffusion models. Two key challenges are addressed by mmDiff: (1) miss-detection of parts of human bodies, which is addressed by a module that isolates feature extraction from different body parts, and (2) signal inconsistency due to environmental interference, which is tackled by incorporating prior knowledge of body structure and motion. Several modules are designed to achieve these goals, whose features work as the conditions for the subsequent diffusion model, eliminating the miss-detection and instability of HPE based on RF-vision. Extensive experiments demonstrate that mmDiff outperforms existing methods significantly, achieving state-of-the-art performances on public datasets.  [^1] The project page of mmDiff is https://fanjunqiao.github.io/mmDiff-site/.
@@ 1 Introduction
Overcoming limitations of camera-based HPE, Radio-Frequency vision (RF-vision) has attracted surging attention for human sensing. The emerging mmWave radar technology presents a promising and feasible solution for HPE due to its price, portability, and energy efficiency [40]. Operated at the frequencies of 30-300 GHz [46], commercial mmWave radars transmit and receive RF signals that penetrate human targets and occlusions. Radar point clouds (PCs) are further extracted as the salient target detection via monitoring signals’ characteristic changes [30, 43]. Therefore, the extracted PCs are more robust to adverse environments [48] and reveal little privacy [44], inspiring accurate and privacy-preserving HPE [22, 44, 2, 43]. However, due to the bandwidth and hardware limitations [26], radar PCs are sparse with limited geometric information [2], leading to huge difficulties in achieving HPE. The sparse PCs are noisy in two aspects: (1) mmWave radar has a lower spatial resolution, leading to PC dispersion throughout the target area accompanied by ghost points caused by multi-path effect [37, 26]; (2) signal’s specular reflection and interference [6, 5] further cause inconsistent sensing data, leading to occasional miss-detection of human parts.
Figure 1: Left: challenges of mmWave PCs. Right: the performance of existing SOTA (P4Transformer [12]) compared to ours. The GTs are black and predictions are colored. PC’s sparsity and dispersion cause inaccurate spline and shoulder. Inconsistent PCs with occasional miss-detection further cause size variance and pose vibration. mmDiff proposes diffusion-based pose estimation with enhanced accuracy and stability.
@@ 3 Methodology
Figure 2: mmDiff proposes a diffusion-based HPE model, using mmWave radar information as conditions. $k\in[0..K]$ denotes the diffusion step. Four modules are proposed as the more reliable guidance, addressing PCs’ noise and inconsistency: GRC and LRC first extract robust global-local radar features, $C^{glo}$ and $C^{loc}$; SLC and TMC then extract consistent human structure and motion patterns, $C^{tem}$ and $C^{lim}$.
We address 3D HPE task with noisy and sparse PCs from mmWave radar. At each time frame $t\geq 0$, given the input 6-dimensional radar point cloud $R_{t}\in\mathbb{R}^{N\times 6}$, where $N$ denotes the number of detected points, our task is to estimate the ground-truth 17-joint human pose $H_{t}=\{h_{t}^{1},h_{t}^{2},...,h_{t}^{17}\}\in\mathbb{R}^{17\times 3}$. Apart from the Cartesian coordinate $\{x,y,z\}$, each detected radar point contains three attributes $\{v,E,A\}$, representing velocity, energy, and amplitude respectively.
@@ 3.1 Diffusion-based Human Pose Estimation
Diffusion models [15, 35] are a category of probabilistic generative models popular in various tasks, e.g., image generation. Given a noisy image from a noisy distribution, diffusion models can generate realistic image samples that match the natural image distribution, through iterative noise removal [23, 18]. Extended to HPE, diffusion models can estimate the distribution of reasonable human poses for realistic pose generation. Particularly with noisy radar modality, miss-detected body joints can be accurately estimated by inferring from the detected ones, and inaccurate joints causing twisted human structures are potentially refined.
@@ 3.2 Global-local Radar Context
Existing mmWave HPE paradigm applies encoder-decoder architectures to encode the radar PCs, $R\in\mathbb{R}^{N\times 6}$, into latent feature representations and decode them into human poses. Anchor-based methods [12, 47] are common options for PC encoders, where point anchors are first sampled from the Farthest Point Sampling algorithm and nearby PCs are extracted as the anchor features. As a result, the holistic radar PCs are encoded into PC features $F^{r}\in\mathbb{R}^{P\times 1024}$, where $P$ indicates the anchor number. To decode the PC features, a Multi-Layer Perceptron (MLP) is commonly applied as a dimension-reducing projection to generate the joint feature $F^{j}\in\mathbb{R}^{17\times 1024}$, which is further decoded into human poses $\tilde{H}\in\mathbb{R}^{17\times 3}$ by another MLP. However, occasional radar miss-detection causes uncertainty within the PC features $F^{r}$, while MLP-based projection can hardly identify miss-detected joints. Additionally, radar low-resolution PCs further impose noises. As a result, the estimated human joints are generally coarse.
@@
LRC further performs local point-to-point self-attention to extract higher-resolution features from local PCs near body joints. To select local PCs, existing methods [43] utilize static joint anchors from coarsely-estimated human poses $\tilde{H}$. However, the local joint features should dynamically reflect the joints’ errors at different diffusion steps. Therefore, we propose dynamic joint anchors from intermediate diffusion poses $\hat{H}_{k}$ for PC selection. With $i\in[1,..,17]$ and each joint $\hat{h}_{k}^{i}$ from the $\hat{H}_{k}$ as an anchor, the $\bar{K}$-nearest-neighbors ($\bar{K}$) algorithm is applied to select $\bar{K}$ nearest points as local PCs, $\bar{R}^{i}_{k}\in\mathbb{R}^{\bar{K}\times 6}$. Each $\bar{R}^{i}_{k}$ is first encoded by a shared MLP $g^{l}$ into a $\mathbb{R}^{\bar{K}\times 64}$ embedding, and then fed into a shared small-scale local transformer $\Phi^{l}$ for point-to-point self-attention. Finally, average pooling is performed to generate $\mathbb{R}^{64}$ embeddings, which are further aggregated (concatenated) for every joint anchor $\tilde{h}^{i}$ as the conditional embedding:
@@ 3.4 Overall Learning Objective
In phase two, the remaining three conditional modules and the diffusion model are trained together, with the phase one parameters frozen. The extracted $F^{j}$ and $\tilde{H}$ serve as the input of structural-motion consistency modules, and $\tilde{H}$ initialize $\hat{H}_{K}$ for the reverse diffusion process. The learning objective is minimizing $\mathcal{L}_{\text{diff}}$, which is the diffusion learning objective following DDPM [15]. To ensure accurate limb-length estimation, an $\mathcal{L}_{1}$ limb regression loss is further designed and integrated:
@@ 4.1 Experiment Setup
mmBody [44] studies the robustness of human sensing with various sensors: RGB, Depth, and mmWave radar. Human skeletons are annotated by the MoCap system. Models are set to train on data collected from 2 standard scenes (Lab1 and Lab2), and tested on 3 basic scenes and 4 adverse scenes (including unseen subjects). Meanwhile, mm-Fi [44] is a larger-scale HPE dataset using a lower-bandwidth mmWave radar with sparser PCs. The skeleton annotations are rather unstable compared to the MoCap annotations, as obtained by RGB using the pretrained HRNet-w48 [36]. Three data-splitting methods are used with a train-test split ratio of 4:1: random, cross-subject, and cross-environment.
@@
Our methods are trained for 100 epochs with a batch size of 1024. The Adam algorithm [19] is used for optimization, with the learning rate as $52e-$ 5, the gradient clip as 1.0, and the momentum as 0.9. The forward/reverse diffusion process is set as $K=25$ steps with a constant $\beta$ sampling of 0.001. An average of 5 hypotheses is recorded for a fair comparison with non-diffusion methods. We choose Point4D [12] as the GRC’s PC encoder for mmBody and PointTransformer [47] for mm-Fi. We set $\bar{K}=50$ for $\bar{K}$ algorithm for LRC, but due to insufficient radar points ($N<100$), the LRC module is neglected for mm-Fi. We further set $\Delta t=8$ for TMC and $\lambda=5$ for SLC. The hyper-parameters are obtained empirically, more precise hyper-parameter tuning tricks such as the Bayesian optimization could lead to better results.
@@
PointTransformer [47] $73.09_{\pm 2.70}$ $55.60_{\pm 1.40}$ $75.96_{\pm 6.90}$ $58.70_{\pm 4.30}$ $88.28_{\pm 4.50}$ $68.79_{\pm 2.79}$
Diffpose* [14] $73.44_{\pm 0.29}$ $56.83_{\pm 0.25}$ $70.31_{\pm 0.27}$ $54.12_{\pm 0.31}$ $86.35_{\pm 0.06}$ $66.87_{\pm 0.17}$
mmDiff(G)* $68.62_{\pm 0.06}$ $53.11_{\pm 0.05}$ $68.46_{\pm 0.06}$ $52.55_{\pm 0.05}$ $85.63_{\pm 0.53}$ $66.43_{\pm 0.28}$
mmDiff(G,S)* $65.72_{\pm 0.08}$ $50.72_{\pm 0.01}$ $67.18_{\pm 0.18}$ $51.85_{\pm 0.05}$ $83.39_{\pm 0.17}$ $64.61_{\pm 0.40}$
mmDiff(G,S,T)* $\textbf{65.26}_{\pm\textbf{0.11}}$ $\textbf{50.35}_{\pm\textbf{0.09}}$ $\textbf{65.62}_{\pm\textbf{0.24}}$ $\textbf{50.23}_{\pm\textbf{0.24}}$ $\textbf{82.73}_{\pm\textbf{0.62}}$ $\textbf{63.87}_{\pm\textbf{0.26}}$
4.1.3 Compared Methods.
(1) RGB, Depth, and RGB(D) [6] are benchmarks using different modalities. (2) mmWave methods: P4Transformer [12] is the benchmark on mmBody containing Point4D Convolution as the PC encoder and a transformer for self-attention; PointTransformer [47] is the benchmark on mm-Fi designed to handle the sparser PCs. mmMesh [43] is implemented as an extra mmWave HPE baseline on mmBody. (3) SOTA camera-based HPE methods: camera-based 3D HPE methods generally perform pose lifting from 2D poses to 3D poses. To modify them to perform radar-based HPE, we train the models to perform pose refinements, from coarse poses $\tilde{H}$ to clean 3D poses. PoseFormer [50] is the transformer-based method with temporal-spatial attention. DiffPose [14] is the diffusion-based method using SOTA graph-based GraFormer [49] as backbones.
4.1.4 Evaluation Metric.
Two evaluation metrics are adopted following [17]: (1) Mean Per Joint Position Error (MPJPE (mm)): the average joint error between ground truth and prediction (after pelvis alignment); and (2) Procrustes Analysis MPJPE (PA-MPJPE (mm)): procrustes methods (translation, rotation, and scaling) are performed before error calculation.
Figure 3: Qualitative visualization of the estimated poses on mmBody dataset. mmDiff demonstrates higher keypoint accuracy. The GTs are black and predictions are colored.
Table 3: Ablation studies of proposed modules on mmBody.
@@ 4.2 Overall Result
Figure 4: (a) shows the pose motion stability on mm-Fi by plotting 5 consecutive frames of poses. mmDiff shows more consistent motion patterns (zoom in for details). (b) shows the motion energy levels on mmBody, where lower AKV indicates better stability.
@@ 4.3 Analytics
In Figure 4 (a), we observe a mixing of incorrect skeletons within the correct skeleton timeline with PointTrans. [47], e.g. chest-expanding and hand-waving. With mmDiff, smooth motion patterns are ensured based on human behavioral prior knowledge. Additionally, our proposed mmDiff can mitigate the uncertain locations of legs and arms (caused by low resolution), demonstrating enhanced pose accuracy. Furthermore, for throwing and kicking actions, our method is more stable compared to RGB-extracted ground truth, as camera-based HPE suffers from self-occlusion. In Figure 4 (b), we further apply the Average Keypoint Velocity (AKV) to quantitatively measure the pose stability. AKV is defined as $E_{i=[0..17]}(||J_{t}^{i}-J^{i}_{t-1}||^{2})$, which measures the average inter-frame joint moving distance, indicating the motion energy level and pose stability. The proposed mmDiff demonstrates enhanced pose stability by minimizing joint vibration and avoiding abrupt pose changes.
@@
To validate the effectiveness of the limb-length consistency, we compare mmDiff w/ and w/o the module ($C^{lim}$) and plot the histograms that indicate the limb-length distribution in Figure 5. The ground truth limb length remains constant for all limbs. We observe both reduced limb-length error (in the estimated arms, legs, and spline) and variance (in forearms and lower legs), indicating enhanced pose accuracy and structural stability. We argue that more accurate limb-length can lead to more accurate pose estimation. Qualitatively in figure 4, the variant height and arm’s length for chest expansion and hand raising are mitigated with mmDiff. However, as in Table 1, SLC has limitations when the sensing distance is long (e.g., in Lab 2) due to harder limb-length estimation.
@@
We provide the efficiency analysis in Table 5, where our proposed modules require substantially little computational resources during phase-two diffusion training and inference. All designed modules have a latency of less than 2ms, demonstrating outstanding efficiency in supporting the iterative diffusion process. Meanwhile, the model’s parameters of our designed modules are relatively small compared to the P4Transformer [12], and require minimum computation complexity, as reflected by GFLOPs. It demonstrates the potential for applications like robotics, the Internet of Things, and edge computing.